### v3

# Convergence on the Proton Drip-Line in Thulium

Abstract
Direct observation of proton emission for very small Q-values is often unfeasible due to the long partial half-lives of the proton emission channel associated with tunneling through the Coulomb barrier. Therefore, proton emitters with very small decay energies may require the masses of both parent and daughter nuclei in order to establish them as proton unbound. Nuclear mass models have been used to predict the proton drip-line of the thulium (Tm) isotopic chain (), but until now the proton separation energy has not been experimentally tested. Mass measurements were performed using a Multiple Reflection Time-Of-Flight Mass Spectrometer (MR-TOF-MS) at TRIUMF’s TITAN facility to conclusively map the limit of proton-bound Tm. The masses of neutron-deficient, 149Tm and 150Tm, combined with measurements of 149m,gEr (which were found to deviate from literature by 150 keV), provide the first experimental confirmation that 149Tm is the first proton-unbound nuclide in the Tm chain. Our measurements also enable the strength of the neutron shell gap to be determined at the Tm proton drip-line, providing evidence supporting its continued existence.
I Introduction
When isotopes become sufficiently neutron-deficient it eventually becomes energetically possible for one or more protons to spontaneously escape the nucleus. This transition within an isotopic chain to one or more protons being unbound is known as the proton drip-line, and its location is of fundamental interest. It can be experimentally determined through measurements of the one-proton separation energy, , and a nucleus is said to lie beyond the proton drip-line when (i.e. it is proton-unbound) [1, 2]. Studying the location of the proton drip-line provides a valuable benchmark for the various nuclear models which predict the properties of unstable nuclei.
Near the proton drip-line, a proton must tunnel through the Coulomb barrier in order to be emitted (i.e. a single-nucleon analogue to -decay). In contrast to alpha decay however, it is simpler to calculate the proton emission decay rate or spectroscopic factor, since the decay does not require the formation of a cluster of nucleons [1]. Proton emission is significantly influenced by the centrifugal barrier and the proton emission half-life at a given proton number, , depends strongly on underlying nuclear structure including the angular momentum carried away by the emitted proton [1][3][4], and the decay energy, [4].
Furthermore, the nuclear shell structure can change towards the drip-line; effects such as the weakening or disappearance of shells (quenching) at the classical magic numbers, and the appearance of new magic numbers have been both predicted and observed [5, 6]. Nuclear masses and binding energies provide indispensable information when studying the limits of nuclear existence and shell structure. They allow calculations of the particle separation energies and can be used to make predictions regarding the underlying nuclear structure and nuclear properties such as half-life, angular momentum, or spectroscopic factors in the case of proton emission.
The mass values of nuclides across the shell closure, formed between the (or nearby or ) and the orbitals, have been intensively studied on the neutron-rich side of the nuclear chart, see e.g. [7, 8, 9, 10, 11, 12]. Corresponding investigations at the neutron-deficient shell closure up to the elements of holmium () and erbium () have been performed by ISOLTRAP [13], the ESR using Schottky Mass Spectrometry [14], and recently in ytterbium () at TITAN [15]. The proton drip-line around is associated with an abrupt change in the deformation [16, 1, 17]. Although masses of neutron-deficient thulium isotopes have been measured by SHIPTRAP [18], measurements remain elusive across the shell closure where the one-proton drip-line is predicted to lie. The MR-TOF-MS technique [19] is well suited for measurements of short-lived isotopes. Here we report the first high-precision mass measurements of neutron-deficient thulium isotopes at the proton drip-line, enabling S2N to be determined across .
I.1 Background
Ground state proton emission can occur for any nucleus beyond the proton drip-line, and was first observed in 1982 in 151Lu and 147Tm [20, 21]. Since then, at least 27 heavy ground state proton emitters have been identified, along with a number of isomeric states [22][23].
While 147Tm has long been known to be a proton emitter, and prior direct mass measurements of 147Tm and 148Tm have been able to constrain the location of the proton drip-line to lie somewhere between and [18], the precise isotope at which Tm becomes proton-unbound remains uncertain due to the thus far unmeasured masses of 149Tm and 150Tm. In cases where neither observation of the decay nor direct mass measurements can be used to establish the drip-line, various theoretical approaches are employed to predict its location, but these models are known to struggle to predict far from stability [24].
Density functional theory and phenomenological descriptions of the nuclear binding energy have been utilized to predict the location of the proton drip-line. For instance, the Skyrme results of [25] can be used to interpolate the of odd-Z nuclei. However, for 149Tm, the choice of energy density functional – such as UNEDF1 [26] or SV-min [27] – leads to competing predictions regarding its position relative to the drip-line. Furthermore, the newer BSkG3 model [28] predicts the drip-line to be even closer to stability.
Bayesian model averaging can be used to assign weights to models in an attempt to improve estimates of the binding energies of heavy nuclei [29]. A recent Bayesian statistical analysis was performed to weight the proton drip-line predictions provided by several different energy density functionals, as well as FRDM2012 and HFB-24 [30]. This work predicted the last proton-bound isotope of Tm to occur at (), but this has never been experimentally verified.The precise location of the thulium proton drip-line remains a significant gap in our experimental knowledge of the heavy, proton-deficient nuclei. The present measurement aims to determine the exact position of the ground state proton drip-line for Tm ().
II Experimental Description
The isotopes under study were produced throughout two experimental campaigns at TRIUMF’s Isotope Separator and Accelerator (ISAC) facility [31] using the ISOL (Isotope Separation On-Line) technique. A 480 MeV proton beam was impinged on a Ta target to produce rare isotopes via spallation reactions and the produced Tm and Er isotopes were surface ionized. The continuous beam of neutron-deficient isotopes was purified using the ISAC high-resolution mass separator [32] (resolving power 2500) to select beams from to before preparation in the TITAN Radio-Frequency Quadrupole (RFQ) cooler-buncher trap [33] for precision mass measurements. The resulting bunches of ions were extracted and their kinetic energy matched for injection into TITAN’s MR-TOF-MS [34][35][36] to obtain a mass spectrum. Ions were again cooled and narrowly bunched in the injection trap of the MR-TOF-MS, before they underwent isochronous (energy independent) reflections between two electrostatic ion mirrors before time-of-flight detection using a MagneTOFTM ion detector manufactured by ETP ion detect. All ions observed in the spectrum were delivered and measured in a 1+ charge state and any possible non-isobars present due to charge exchange were removed using a Mass Range Selector [37, 38]. An example spectrum of is given in Figure 1.
Measurements of the nuclei furthest from stability were made possible by employing so-called mass selective re-trapping [39, 40], during which the ion bunch is captured a second time in the injection trap after a first stage of mass separation in the MR-TOF-MS. Mass selective re-trapping can be used to selectively reduce the intensity of the dominant contaminant species by more than three orders of magnitude [15]. This improved the signal-to-background ratio of the 149,150Tm measurements and allowed the MR-TOF-MS to accept a higher overall beam rate while keeping the total number of ions detected per mass measurement cycle within an acceptable range.
During the initial experiment (experiment #1), isotopes were produced using a high-power tantalum target which is equipped with greater heat dissipation, whereas the second experiment (experiment #2) utilized a low-power tantalum target, for which an improved and more homogeneous thermal profile is believed to have enhanced the release of Tm and Er. In both experiments the proton beam from TRIUMF’s main driver cyclotron was impinged on the Ta production targets with beam currents ranging from approximately 25 to 50 A.
The mass of 150Tm was measured along with the Yb masses [15] obtained in experiment #1, which additionally provided anchor masses for alpha decay chains [41]. The resolving power of 270,000 in this experiment resulted in a proton separation energy of 150Tm which was very close to zero, and no definitive conclusion could be drawn as to whether this nucleus is proton-bound. The Tm data-set was greatly improved during experiment #2 by performing additional measurements which extended data to 149Tm, and improved on the initial experimental uncertainties. In particular, experiment #2 benefited from improvements that were made to the MR-TOF-MS system to achieve 400,000 resolving power and lower systematic uncertainty.
II.1 Analysis Procedure
The analysis procedure followed the description in [42]. At all mass units, we utilized a high abundance peak to correct for time-dependent drifts in the Time of Flight (TOF) spectrum using in-house developed data acquisition software [35]. Subsequently, masses were determined by fitting the peaks in each spectrum using a hyper-exponentially-modified Gaussian (hyperEMG) lineshape [43] and the emgfit python library available at [44]. This procedure involves first fitting an isolated high-statistics peak with a multi-dimensional function to precisely determine the shape (i.e. functional form) of a single peak (the shape calibrant) and then using multiple copies of this shape to fit all peaks in the spectrum. Next, the masses of all peaks in the spectrum were calibrated against a single species with a well-established mass (the mass calibrant). Appropriate shape and mass calibrant peaks were chosen for each set of isobars measured to minimize the influence of overlapping isobars and nuclear isomers. In the present analysis, the peaks corresponding to 149Dy, 150Ho, 151Ho, 152Dy, 153Dy, 154Dy, 155Yb, 156Yb, and 157Yb were selected as shape calibrants in their respective isobaric mass spectra. The chosen mass calibrant at each mass number was cross-checked for consistency with the masses of several different isobars listed in the 2020 Atomic Mass Evaluation (AME2020) to confirm its identity, and are listed in Table 1.
II.2 Results
In this work, mass measurements across neutron-deficient Tm and Er isotopes were obtained, with 149Tm, 150Tm representing first-time measurements. Mass excess values are defined as the sum of the ionic mass and the electron mass (since all ions were singly-charged) minus the mass number .
The final mass values were obtained by adding the experimental and known systematic uncertainties in quadrature and making corrections for any known or expected unresolved isomers. These are summarized in Table 1. Additionally, Table 2 contains the uncorrected measurement data – excluding any isomer corrections – and is provided as a reference should new information about these isomers become available in the future. Mass excess (ME) uncertainties include the uncertainty of the calibrant mass, while the Ion Mass Ratio uncertainties do not. The systematic uncertainty was dominated by nonideal switching of voltages in the ion mirror on the extraction side, and was measured using stable ions as described in [45] and [42]. The single measurement presented here from experiment #1, 150Tm, includes a systematic uncertainty of , while the measurements obtained during experiment #2 incorporate a systematic uncertainty of (except A=153). This reduction was the result of improvements made to the tune of the MR-TOF-MS and the stability of the high-voltage electrodes.
The crowded spectrum shown in Figure 2 was the most challenging from which to obtain a high-statistics shape calibrant in the analysis. An isobaric species lies in each of the two tails of the 153Dy shape/mass calibrant, and the functional form needed to be manually selected in order to ignore excessive influence on the fitting algorithm from the isobar in the lower mass tail. Furthermore, the mass 153 spectrum had a total rate of 3.8 ions per trapping cycle, with 2.7 per cycle in the 153Dy calibrant peak. Therefore, an additional systematic uncertainty of per ion per cycle ( total) was required to account for ion-ion interactions (based on off-line tests).
According to ENSDF, 153Tm has a long-lived isomer at 43.20.2 keV [46] and 155Tm one at 416 keV [47]. 151Tm, 152Tm, and 154Tm respectively contain a s, a s, and a s isomer [48, 49, 50] for which the excitation energy was taken from NUBASE 2020 [51] to be 936 keV, -100250 keV, and 7050 keV. NUBASE 2020 also infers excitation energies for 149Tm (100#50# keV) [52] and 150Tm (140#140# keV) [53] from trends in neighbouring nuclei; the higher 671 keV, 5.2 ms isomer of 150Tm is not expected to be observable in this experiment due to its short half-life and no indication of it was observed. 157Lu also contains an unknown contribution from a 20.92.0 keV isomer (4.8 s) [54].
Since these isomeric contributions may shift the determined mass to appear heavier than for a peak containing only the ground state, the standard procedure outlined in Appendix B of AME2020 [55] was used to determine an appropriate correction to the mass and the corresponding increase in uncertainty. The ground state mass () was obtained from the experimental mass () and the previously reported excitation energy () through the relation , and additional uncertainties of , and were added in quadrature with the experimental error, where is the uncertainty in the reported excitation energy. The Mass Excesses and Ion Mass Ratio was thus corrected for all species likely containing an unresolved isomer (e.g. 13 keV for 153gTm and 12 keV for 155gTm).
The mass spectra of 149,150Tm strongly benefited from using mass selective re-trapping to improve the Tm-to-contaminant ratio by about three to four orders of magnitude depending on their separation, as shown in Figure 1 of reference [15]. Once applied, the peaks were observable above the background and could be fit.
As seen in Figure 3, our measured mass values show overall agreement with the literature values from AME2020, including for 157Tm and 156Tm which are not known to have long-lived isomers or isobaric species that could impact our measurements. The two new mass values of 149,150Tm now close a gap in the long chain of previously measured neutron-deficient isotopes of Tm [18, 56, 57], with mass data now spanning from the proton emitter 147Tm at N=78 to stable 169Tm at N=100. Alongside the measured Tm isotopes, 149-154Er were also identified in the mass spectra. The masses of 150-154Er all lie within 1.5 of the presently accepted values in the AME2020 and any known isomers are MeV above the ground state and thus do not impact the measurements. After correcting for the isomer, our measurement of 157Lu yields a mass 5629 keV heavier than the AME value.
149gEr and 149mEr (bottom Figure 1) were fit using two hyperEMG functions calibrated to the shape of the 149Dy peak. The known isomeric state was delivered to TITAN at greater intensity and the ground state peak shows up as an excess in the left shoulder. Our 149gEr and 149mEr masses were found to be heavier by 2.9 and 3.9 respectively when compared to the literature value obtained via a measurement of 149mEr at the GSI ESR [14]. This results in a shift of the of 150Tm (see below).
| Species | Calibrant | Ion Mass Ratio | ME (keV) | ME (keV) | TITAN - AME (keV) |
| 157gLu † | 157Yb | 1.00004813(16) | -46384 (26) | -46440 (12) | 56(29) |
| 157Tm | 157Yb | 0.99996379(18) | -58714 (28) | -58709 (28) | -4(40) |
| 156Tm | 156Yb | 0.99997554(15) | -56818 (24) | -56834 (14) | 17(28) |
| 155gTm † | 155Er | 1.00003875(17) | -56617 (26) | -56626 (10) | 9(28) |
| 155Yb | 155Er | 1.00008157(15) | -50437 (23) | -50503 (17) | 66(28) |
| 154gTm * | 154Dy | 1.00011154(29) | -54401 (42) | -54427 (14) | 26(44) |
| 154Er | 154Dy | 1.00005458(16) | -62569 (24) | -62605 (5) | 36(24) |
| 153gTm † | 153Dy | 1.00010661(37) | -53957 (54) | -53973 (12) | 16(55) |
| 153Er | 153Dy | 1.00006064(34) | -60505 (48) | -60467 (9) | -39(49) |
| 152gTm * | 152Dy | 1.00013131(94) | -51535 (133) | -51720 (50) | 186(142) |
| 152Er | 152Dy | 1.00006789(18) | -60510 (26) | -60500 (9) | -10(27) |
| 151gTm * | 151Dy | 1.00012788(26) | -50774 (36) | -50772 (19) | -3(41) |
| 151Er | 151Dy | 1.00007472(15) | -58248 (22) | -58266 (16) | 19(27) |
| 150gTm ** | 150Dy | 1.00016335(62) | -46497 (87) | -46490 (200#) | ## |
| 150Er | 150Dy | 1.00008237(16) | -57807 (23) | -57831 (17) | 24(29) |
| 149gTm ** | 149Dy | 1.00017215(36) | -43812 (51) | -43940 (200#) | ## |
| 149gEr | 149Dy | 1.00010171(33) | -53584 (47) | -53742 (28) | 158(55) |
| 149mEr | 149Dy | 1.00010695(16) | -52857 (25) | -53000 (28) | 143(37) |
| 150gTm (Exp 1)** | 150Dy | 1.00016357(73) | -46467 (103) | -46490 (200#) | ## |
III Discussion
As detailed below, our measurements of the Tm and Er mass chains establish 149Tm as the first proton-unbound Tm isotope. Furthermore, they support recent evidence for the persistence of the shell closure at [15] with a corresponding measurement of the trend in at .
Last Proton-Bound Isotope Determined from 149m,149gEr and 150gTm
The Er isotopic chain was measured alongside Tm during experiment #2 from 154Er to 149Er. During this second measurement, a discrepancy in the mass of 149gEr relative to the literature value was found which significantly affects the one-proton separation energy () of the 150Tm ground state. Together, the newly-measured 150Tm mass and the AME2020 mass of 149gEr results in a that is well within of zero, but using both updated masses yields a positive proton separation energy consistent with proton-bound 150Tm ().
As with 151Yb, for which the ISAC target was seen to preferentially deliver the spin 11/2 isomer over the ground state, 149Er, having two fewer protons, similarly appeared to favour the excited state. We therefore fit the two-component 149Er peak using two hyperEMG functions having the same parameters as the 149Dy peak; this yields mass values for both the dominant isomeric state (0.2 s), as well as the ground state (2 s) which shows up as an excess within the low-mass tail of the isomer peak.
The ground state mass of 149Er can be determined by two different methods using our data. One approach is to extract it from the measured isomer mass and the known excitation energy as was previously done in an experiment using Schottky mass spectrometry [14], which reported an uncertainty of 28 keV. The other method is to obtain the mass directly from the fit of the ground state within the tail of the isomer using the known peak shape.
The first approach obtains the mass of 149gEr using the known isomeric excitation energy which was initially measured via a two -ray internal transition to the ground state proceeding through a 630.5 keV (M4) 111.0 keV (M1) cascade [58] This was later confirmed in subsequent experiments which yielded 630.3 keV 111.3 keV [59] and 630.52.6 keV 111.31.1 keV by [60]. The 149mEr isomer is established to lie 741.80.2 keV above the ground state according to the NUBASE 2020 evaluation [51].
This excitation energy is consistent with our recent result in which the newly measured masses of 151g+mYb were used to support a theoretical explanation for the remarkably consistent first excitation energy (approximately 750 keV) in the isotonic chain containing 149Er [15]. This has been attributed to a clustering of single-proton energy levels for the mildly deformed nuclei at [15] resulting from the single hole in the neutron orbital common to 149Er and many of its isotones [58]. This provides increased confidence in the reliability of the established excitation energy of the long-lived 741.690.23 keV isomer 149mEr (9.60.6 s). Using this value with our current mass measurement of 149mEr to indirectly determine the ground state mass yields a 149gEr mass excess of -5359925 keV, a 143 keV increase from AME2020.
However, the present dataset also enables a more direct measurement of the ground state mass by utilizing the lower abundance, overlapping TOF peak of the ground state of 149Er and employing a hyperEMG fitting procedure. Directly fitting the lower abundance 149gEr (as in Figure 1) with the hyperEMG function calibrated to the shape of the 149Dy peak yields a mass excess of -5358447 keV, an increase in mass of 15855 keV from the AME2020 value. The excitation energy of 149mEr can then be determined from the difference between our measurements of ground state and isomer to be 72653 keV.
While the newly determined mass of 149gEr and the isomer 149mEr deviate from the mass value in the AME2020, they are consistent with observable trends in the binding energy of nuclides in this region. The cause of our discrepancy in our measurement of this single mass is unknown, but may warrant a third measurement of the 149g,mEr isotope for confirmation.
These new masses of the 149Er isomer and ground state lead us to adopt a revised one-proton separation energy of the ground state which directly impacts the location of the proton drip-line in Tm, since the proton separation energy of a nucleus depends on both the mass of the candidate proton emitter as well as on the mass of the daughter nucleus of the decay. Both the 149,150Tm masses as well as the revised 149Er value were thus needed for an unambiguous identification of the Tm drip-line, since the new Tm masses alone did not exclude the possibility that 150Tm could also be proton-unbound when using the 149Er value from the AME2020. These measurements confirm that 150Tm is indeed proton-bound, agreeing qualitatively with the Bayesian model averaging predictions of [30] that 149Tm is the first proton-unbound thulium isotope. Furthermore, the proton emission Q-value of 149Tm has been determined and can now be used as an input for decay rate calculations.
149Tm: The First p-Unbound Isotope
The proton emission energy can be determined from . Our direct measurements of 149gEr and 150gTm show 150gTm to be proton-bound, having a Q-value of -20299 keV. Meanwhile, our measurement of 149gTm establishes it as the first proton-unbound isotope in the thulium chain. For it we obtain a proton emission Q-value of +37852 keV when calculated using the measured mass of 149gTm and the masses of 148Er and 1H from the AME2020. This decay energy is on the high side of the AME extrapolation of 250201 keV and can be used in estimations of the proton emission half-life for a given angular momentum transfer (e.g. [3, 4]). The small Q-value of the 149gTm decay allows for it to proceed to only one possible final state, the 0+ ground state of 148Er, so it is expected that J. Both of the above decay Q-values and their uncertainties take into account the expected unresolved isomers listed in NUBASE inferred from trends in neighbouring nuclei.
Evolution of the Neutron Shell Closure
In addition to determining the location of the drip-line, the new masses of 149/150Tm allow for a comparison to predictions of the empirical shell gap of the shell closure defined by (see Figure 5). The models shown include UNEDF0 [62], UNEDF1 [26], HFB21 [63], Duflo-Zuker [64], FRDM [65], SkP [66], SV-min [27], and ETFSI-Q [67]. In this region most models over-predict the shell gap of the N=82 isotone chain by 0.5-3 MeV with the exceptions of Duflo-Zuker under-predicting by nearly 1 MeV and HFB21 which lies within 500 keV of the isotones from to .
The present work measures the of Tm at the N=82 shell closure and shows that it is, in fact, the first isotone for which depends on the binding energy of a proton-unbound nucleus (149Tm). These measurements indicate the persistence of the shell closure, which is consistent with the recent corresponding measurement of Yb [15]. Furthermore, the empirical shell gap at Z=69 continues the trend seen in the more stable nuclei such as Z=63 and 67 of having a local maximum in at the odd-Z nuclei.
IV Conclusion and Outlook
The masses of neutron deficient Tm and Er isotopes obtained in this work identify the precise location of the proton drip-line in the Tm isotopic chain. The transition from positive to negative proton separation energy is observed between 149gTm and 150gTm. Our updated mass of 149Er directly impacts the proton-bound nature of 150Tm since it would be the daughter nucleus of the proton emission if this decay mode were possible. The masses of two previously unmeasured Tm isotopes were measured in this experiment. Additonally, an update to the previously measured mass of 149mEr and a first direct measurement of the ground state 149gEr determined them to be 150 keV heavier than the current literature values, thereby establishing 150gTm as proton-bound.
The well-understood composition of the mass spectra provides confidence in our identification of isotopes using the MR-TOF-MS. The mass of 149gEr was determined directly by performing hyperEMG fits on the time-of-flight spectrum obtained with TITAN’s MR-TOF-MS and indirectly through the 149mEr isomer. Together, our masses of 149gTm, 150gTm, and 149gEr establish that 149gTm, having 20 fewer neutrons than the only stable isotope of Tm (169Tm, ), is the first proton-unbound ground state in this isotopic chain. This agrees with the drip-line prediction obtained from a recent Bayesian analysis of various theoretical models [30].
Our update to the literature masses also highlights the importance of measuring and reporting not only the masses of previously unmeasured isotopes but also confirmations of literature mass values which provide input for various models and predictions. Heavier exotic species pose challenges that arise from the abundance of isomers and isobaric molecules that must be identified and, therefore, must be treated with caution. The mass values presented here provide a new benchmark for the models used to predict the binding energies of unmeasured nuclei in the neutron-deficient region near N=82 and establish the Z=69 proton drip-line.
V Acknowledgements
TITAN is funded by the Natural Sciences and Engineering Research Council (NSERC) of Canada and through TRIUMF by the National Research Council (NRC) of Canada.
This work was further supported by the German Research Foundation (DFG), Grant No. 422761894; by the German Federal Ministry for Education and Research (BMBF), Grants No. 05P16RGFN1 and No. 05P19RGFN1; by the Hessian Ministry for Science and Art through the LOEWE Center HIC for FAIR; by the JLU and GSI Helmholtzzentrum für Schwerionenforschung under the JLU-GSI strategic Helmholtz partnership agreement.
For the purpose of open access, the authors have applied a creative commons attribution (CC BY) licence to any author accepted manuscript version arising.
| Species | Calibrant | Ion Mass Ratio | ME (keV) | ME (keV) |
| 157g+mLu | 157Yb | 1.00004820(16) | -46373 (25) | -46440 (12) |
| 157Tm | 157Yb | 0.99996379(18) | -58714 (28) | -58709 (28) |
| 156Tm | 156Yb | 0.99997554(15) | -56818 (24) | -56834 (14) |
| 155g+mTm * | 155Er | 1.00003889(15) | -56596 (23) | -56626 (10) |
| 155Yb | 155Er | 1.00008157(15) | -50437 (23) | -50503 (17) |
| 154g+mTm * | 154Dy | 1.00011179(18) | -54366 (26) | -54427 (14) |
| 154Er | 154Dy | 1.00005458(16) | -62569 (24) | -62605 (5) |
| 153g+mTm * | 153Dy | 1.00010676(36) | -53935 (52) | -53973 (12) |
| 153Er | 153Dy | 1.00006064(34) | -60505 (48) | -60467 (9) |
| 152g+mTm * | 152Dy | 1.00013096(23) | -51585 (33) | -51720 (50) |
| 152Er | 152Dy | 1.00006789(18) | -60510 (26) | -60500 (9) |
| 151g+mTm * | 151Dy | 1.00012821(17) | -50728 (24) | -50772 (19) |
| 151Er | 151Dy | 1.00007472(15) | -58248 (22) | -58266 (16) |
| 150Tm * | 150Dy | 1.00016386(23) | -46427 (33) | -46490 (200#) |
| 150Er | 150Dy | 1.00008237(16) | -57807 (23) | -57831 (17) |
| 149Tm * | 149Dy | 1.00017251(24) | -43762 (34) | -43940 (200#) |
| 149gEr | 149Dy | 1.00010171(33) | -53584 (47) | -53742 (28) |
| 149mEr | 149Dy | 1.00010695(17) | -52857 (25) | -53000 (28) |
| 150Tm (Exp 1) * | 150Dy | 1.00016407(45) | -46397 (63) | -46490 (200#) |

### v5

@@
 Abstract
-Direct observation of proton emission for very small Q-values is often unfeasible due to the long partial half-lives of the proton emission channel associated with tunneling through the Coulomb barrier.
+Direct observation of proton emission for very small ·-values is often unfeasible due to the long partial half-lives of the proton emission channel associated with tunneling through the Coulomb barrier.
 Therefore, proton emitters with very small decay energies may require the masses of both parent and daughter nuclei in order to establish them as proton unbound.
@@
 it is proton-unbound) [1, 2].
-Studying the location of the proton drip-line provides a valuable benchmark for the various nuclear models which predict the properties of unstable nuclei.
+Studying the location of the proton drip-line provides a valuable benchmark for the various nuclear models that predict the properties of unstable nuclei.
 Near the proton drip-line, a proton must tunnel through the Coulomb barrier in order to be emitted (i.e.
 a single-nucleon analogue to -decay).
-In contrast to alpha decay however, it is simpler to calculate the proton emission decay rate or spectroscopic factor, since the decay does not require the formation of a cluster of nucleons [1].
-Proton emission is significantly influenced by the centrifugal barrier and the proton emission half-life at a given proton number, , depends strongly on underlying nuclear structure including the angular momentum carried away by the emitted proton [1][3][4], and the decay energy, [4].
-Furthermore, the nuclear shell structure can change towards the drip-line; effects such as the weakening or disappearance of shells (quenching) at the classical magic numbers, and the appearance of new magic numbers have been both predicted and observed [5, 6].
+In contrast to decay, however, it is simpler to calculate the proton emission decay rate or spectroscopic factor, since the decay does not require the formation of a cluster of nucleons [1].
+Proton emission is significantly influenced by the centrifugal barrier, and the proton emission half-life at a given proton number depends strongly on underlying nuclear structure, including the angular momentum carried away by the emitted proton [1][3][4] and the decay energy [4].
+Furthermore, the nuclear shell structure can change towards the drip-line; effects such as the weakening or disappearance of shells (quenching) at the classical magic numbers and the appearance of new magic numbers have been both predicted and observed [5, 6].
 Nuclear masses and binding energies provide indispensable information when studying the limits of nuclear existence and shell structure.
 They allow calculations of the particle separation energies and can be used to make predictions regarding the underlying nuclear structure and nuclear properties such as half-life, angular momentum, or spectroscopic factors in the case of proton emission.
-The mass values of nuclides across the shell closure, formed between the (or nearby or ) and the orbitals, have been intensively studied on the neutron-rich side of the nuclear chart, see e.g.
-[7, 8, 9, 10, 11, 12].
-Corresponding investigations at the neutron-deficient shell closure up to the elements of holmium () and erbium () have been performed by ISOLTRAP [13], the ESR using Schottky Mass Spectrometry [14], and recently in ytterbium () at TITAN [15].
+The mass values of nuclides across the shell closure, formed between the (or nearby or ) and the orbitals, have been intensively studied on the neutron-rich side of the nuclear chart; see, e.g., [7, 8, 9, 10, 11, 12].
+Corresponding investigations at the neutron-deficient shell closure up to the elements of holmium () and erbium () have been performed at ISOLTRAP (ISOLDE/CERN) using Penning trap [13], at the ESR storage ring (GSI) using Schottky mass spectrometry [14], and recently for ytterbium () at TITAN [15] using the multiple reflection time- of-flight mass spectrometer (MR-TOF-MS) technique [16], which is well suited for measurements of short-lived isotopes.
 The proton drip-line around is associated with an abrupt change in the deformation [17, 1, 18].
 Although masses of neutron-deficient thulium isotopes have been measured by SHIPTRAP [19], measurements remain elusive across the shell closure where the one-proton drip-line is predicted to lie.
-The MR-TOF-MS technique [19] is well suited for measurements of short-lived isotopes.
-Here we report the first high-precision mass measurements of neutron-deficient thulium isotopes at the proton drip-line, enabling S2N to be determined across .
+Using the MR-TOF-MS at TITAN, here we report the first high-precision mass measurements of neutron-deficient thulium isotopes at the proton drip-line, enabling S2N to be determined across .
 I.1 Background
@@
 Ions were again cooled and narrowly bunched in the injection trap of the MR-TOF-MS, before they underwent isochronous (energy independent) reflections between two electrostatic ion mirrors before time-of-flight detection using a MagneTOFTM ion detector manufactured by ETP Ion Detect.
-All ions observed in the spectrum were delivered and measured in a 1+ charge state and any possible non-isobars present due to charge exchange were removed using a Mass Range Selector [37, 38].
-An example spectrum of is given in Figure 1.
-Measurements of the nuclei furthest from stability were made possible by employing so-called mass selective re-trapping [39, 40], during which the ion bunch is captured a second time in the injection trap after a first stage of mass separation in the MR-TOF-MS.
-Mass selective re-trapping can be used to selectively reduce the intensity of the dominant contaminant species by more than three orders of magnitude [15].
+All ions observed in the spectrum were delivered and measured in a 1+ charge state and any possible non-isobars present due to charge exchange were removed using a mass range selector [37].
+An example spectrum of is given in Fig.
+1.
+Measurements of the nuclei furthest from stability were made possible by employing so-called mass selective retrapping [38, 39], during which the ion bunch is captured a second time in the injection trap after a first stage of mass separation in the MR-TOF-MS.
+Mass selective retrapping can be used to selectively reduce the intensity of the dominant contaminant species by more than three orders of magnitude [15].
 This improved the signal-to-background ratio of the 149,150Tm measurements and allowed the MR-TOF-MS to accept a higher overall beam rate while keeping the total number of ions detected per mass measurement cycle within an acceptable range.
-During the initial experiment (experiment #1), isotopes were produced using a high-power tantalum target which is equipped with greater heat dissipation, whereas the second experiment (experiment #2) utilized a low-power tantalum target, for which an improved and more homogeneous thermal profile is believed to have enhanced the release of Tm and Er.
+During the initial experiment (experiment 1), isotopes were produced using a high-power tantalum target which is equipped with greater heat dissipation, whereas the second experiment (experiment 2) utilized a low-power tantalum target, for which an improved and more homogeneous thermal profile is believed to have enhanced the release of Tm and Er.
 In both experiments the proton beam from TRIUMF’s main driver cyclotron was impinged on the Ta production targets with beam currents ranging from approximately 25 to 50 A.
-The mass of 150Tm was measured along with the Yb masses [15] obtained in experiment #1, which additionally provided anchor masses for alpha decay chains [41].
-The resolving power of 270,000 in this experiment resulted in a proton separation energy of 150Tm which was very close to zero, and no definitive conclusion could be drawn as to whether this nucleus is proton-bound.
-The Tm data-set was greatly improved during experiment #2 by performing additional measurements which extended data to 149Tm, and improved on the initial experimental uncertainties.
-In particular, experiment #2 benefited from improvements that were made to the MR-TOF-MS system to achieve 400,000 resolving power and lower systematic uncertainty.
+The mass of 150Tm was measured along with the Yb masses [15] obtained in experiment 1, which additionally provided anchor masses for decay chains [40].
+The resolving power of 270 000 in this experiment resulted in a proton separation energy of 150Tm that was very close to zero, and no definitive conclusion could be drawn as to whether this nucleus is proton-bound.
+The Tm data-set was greatly improved during experiment 2 by performing additional measurements that extended data to 149Tm, and improved on the initial experimental uncertainties.
+In particular, experiment 2 benefited from improvements that were made to the MR-TOF-MS system to achieve 400 000 resolving power and lower systematic uncertainty.
 II.1 Analysis Procedure
 The analysis procedure followed the description in [41].
-At all mass units, we utilized a high abundance peak to correct for time-dependent drifts in the Time of Flight (TOF) spectrum using in-house developed data acquisition software [35].
+At all mass units, we utilized a high abundance peak to correct for time-dependent drifts in the time of flight (TOF) spectrum using in-house-developed data acquisition software [35].
 Subsequently, masses were determined by fitting the peaks in each spectrum using a hyper-exponentially-modified Gaussian (hyperEMG) lineshape [42] and the emgfit Python library available at [43].
-This procedure involves first fitting an isolated high-statistics peak with a multi-dimensional function to precisely determine the shape (i.e.
-functional form) of a single peak (the shape calibrant) and then using multiple copies of this shape to fit all peaks in the spectrum.
+This procedure involves first fitting an isolated high-statistics peak with a multi-dimensional function to precisely determine the shape (i.e., functional form) of a single peak (the shape calibrant) and then using multiple copies of this shape to fit all peaks in the spectrum.
 Next, the masses of all peaks in the spectrum were calibrated against a single species with a well-established mass (the mass calibrant).
@@
 In the present analysis, the peaks corresponding to 149Dy, 150Ho, 151Ho, 152Dy, 153Dy, 154Dy, 155Yb, 156Yb, and 157Yb were selected as shape calibrants in their respective isobaric mass spectra.
-The chosen mass calibrant at each mass number was cross-checked for consistency with the masses of several different isobars listed in the 2020 Atomic Mass Evaluation (AME2020) to confirm its identity, and are listed in Table 1.
+The chosen mass calibrant at each mass number was cross-checked for consistency with the masses of several different isobars listed in the 2020 Atomic Mass Evaluation (AME2020) to confirm its identity, and is listed in Table 1.
 II.2 Results
-In this work, mass measurements across neutron-deficient Tm and Er isotopes were obtained, with 149Tm, 150Tm representing first-time measurements.
+In this work, mass measurements across neutron-deficient Tm and Er isotopes were obtained, with 149Tm and 150Tm representing first-time measurements.
 Mass excess values are defined as the sum of the ionic mass and the electron mass (since all ions were singly-charged) minus the mass number .
@@
 The systematic uncertainty was dominated by nonideal switching of voltages in the ion mirror on the extraction side, and was measured using stable ions as described in [44] and [41].
-The single measurement presented here from experiment #1, 150Tm, includes a systematic uncertainty of , while the measurements obtained during experiment #2 incorporate a systematic uncertainty of (except A=153).
-This reduction was the result of improvements made to the tune of the MR-TOF-MS and the stability of the high-voltage electrodes.
-The crowded spectrum shown in Figure 2 was the most challenging from which to obtain a high-statistics shape calibrant in the analysis.
+The single measurement presented here from experiment 1, 150Tm, includes a systematic uncertainty of , while the measurements obtained during experiment 2 incorporate a systematic uncertainty of (except A=153).
+This reduction was the result of improvements made to the tuning of the MR-TOF-MS and the stability of the high-voltage electrodes.
+The crowded spectrum shown in Fig.
+2 was the most challenging from which to obtain a high-statistics shape calibrant in the analysis.
 An isobaric species lies in each of the two tails of the 153Dy shape/mass calibrant, and the functional form needed to be manually selected in order to ignore excessive influence on the fitting algorithm from the isobar in the lower mass tail.
@@
 Therefore, an additional systematic uncertainty of per ion per cycle ( total) was required to account for ion-ion interactions (based on off-line tests).
-According to ENSDF, 153Tm has a long-lived isomer at 43.20.2 keV [46] and 155Tm one at 416 keV [47].
-151Tm, 152Tm, and 154Tm respectively contain a s, a s, and a s isomer [48, 49, 50] for which the excitation energy was taken from NUBASE 2020 [51] to be 936 keV, -100250 keV, and 7050 keV.
-NUBASE 2020 also infers excitation energies for 149Tm (100#50# keV) [52] and 150Tm (140#140# keV) [53] from trends in neighbouring nuclei; the higher 671 keV, 5.2 ms isomer of 150Tm is not expected to be observable in this experiment due to its short half-life and no indication of it was observed.
+According to ENSDF, 153Tm has a long-lived isomer at 43.20.2 keV [45] and 155Tm one at 416 keV [46].151Tm, 152Tm, and 154Tm respectively contain isomers with half-lives , a , and s [47, 48, 49], for which the excitation energies were taken from NUBASE 2020 [50] to be 936, -100250, and 70 50 keV.
+NUBASE 2020 also lists extrapolated excitation energies for the unmeasured isotopes 149Tm (10050 keV) [51] and 150Tm (140140 keV) [52] determined from trends in neighbouring nuclei; the higher 671 keV, 5.2 ms isomer of 150Tm is not expected to be observable in this experiment due to its short half-life, and no indication of it was observed.
 157Lu also contains an unknown contribution from a 20.92.0 keV isomer (4.8 s) [53].
 Since these isomeric contributions may shift the determined mass to appear heavier than for a peak containing only the ground state, the standard procedure outlined in Appendix B of AME2020 [54] was used to determine an appropriate correction to the mass and the corresponding increase in uncertainty.
-The ground state mass () was obtained from the experimental mass () and the previously reported excitation energy () through the relation , and additional uncertainties of , and were added in quadrature with the experimental error, where is the uncertainty in the reported excitation energy.
-The Mass Excesses and Ion Mass Ratio was thus corrected for all species likely containing an unresolved isomer (e.g.
-13 keV for 153gTm and 12 keV for 155gTm).
-The mass spectra of 149,150Tm strongly benefited from using mass selective re-trapping to improve the Tm-to-contaminant ratio by about three to four orders of magnitude depending on their separation, as shown in Figure 1 of reference [15].
+The ground state mass () was obtained from the experimental mass () and the previously reported excitation energy () through the relation , and additional uncertainties of and were added in quadrature with the experimental error, where is the uncertainty in the reported excitation energy.
+The mass excesses and ion mass ratio were thus corrected for all species likely containing an unresolved isomer (e.g., 13 keV for 153gTm and 12 keV for 155gTm).
+The mass spectra of 149,150Tm strongly benefited from using mass selective retrapping to improve the Tm-to-contaminant ratio by about three to four orders of magnitude depending on their separation, as shown in Fig.
+1 of Ref.
+[15].
 Once applied, the peaks were observable above the background and could be fit.
-As seen in Figure 3, our measured mass values show overall agreement with the literature values from AME2020, including for 157Tm and 156Tm which are not known to have long-lived isomers or isobaric species that could impact our measurements.
-The two new mass values of 149,150Tm now close a gap in the long chain of previously measured neutron-deficient isotopes of Tm [18, 56, 57], with mass data now spanning from the proton emitter 147Tm at N=78 to stable 169Tm at N=100.
+As seen in Figure 3, mass values show overall agreement with the literature values from AME2020, including those for 157Tm and 156Tm, which are not known to have long-lived isomers or isobaric species that could here impact them.
+The two new mass values of 149,150Tm now close a gap in the long chain of previously measured neutron-deficient isotopes of Tm [19, 55, 56], with mass data now spanning from the proton emitter 147Tm at to stable 169Tm at .
 Alongside the measured Tm isotopes, 149-154Er were also identified in the mass spectra.
-The masses of 150-154Er all lie within 1.5 of the presently accepted values in the AME2020 and any known isomers are MeV above the ground state and thus do not impact the measurements.
+The masses of 150-154Er all lie within 1.5 of the presently accepted values in the AME2020, and any known isomers are MeV above the ground state and thus do not impact the measurements.
 After correcting for the isomer, our measurement of 157Lu yields a mass 5629 keV heavier than the AME value.
-149gEr and 149mEr (bottom Figure 1) were fit using two hyperEMG functions calibrated to the shape of the 149Dy peak.
+149gEr and 149mEr (bottom Fig.
+1) were fit using two hyperEMG functions calibrated to the shape of the 149Dy peak.
 The known isomeric state was delivered to TITAN at greater intensity and the ground state peak shows up as an excess in the left shoulder.
@@
 This results in a shift of the of 150Tm (see below).
-| Species | Calibrant | Ion Mass Ratio | ME (keV) | ME (keV) | TITAN - AME (keV) |
+| Species | Calibrant | Ion Mass Ratio | ME (keV) | ME (keV) | TITAN – AME (keV) |
 | 157gLu † | 157Yb | 1.00004813(16) | -46384 (26) | -46440 (12) | 56(29) |
@@
 | 149mEr | 149Dy | 1.00010695(16) | -52857 (25) | -53000 (28) | 143(37) |
-| 150gTm (Exp 1)** | 150Dy | 1.00016357(73) | -46467 (103) | -46490 (200#) | ## |
+| 150gTm (Expt. 1)** | 150Dy | 1.00016357(73) | -46467 (103) | -46490 (200#) | ## |
 III Discussion
@@
 Last Proton-Bound Isotope Determined from 149m,149gEr and 150gTm
-The Er isotopic chain was measured alongside Tm during experiment #2 from 154Er to 149Er.
-During this second measurement, a discrepancy in the mass of 149gEr relative to the literature value was found which significantly affects the one-proton separation energy () of the 150Tm ground state.
-Together, the newly-measured 150Tm mass and the AME2020 mass of 149gEr results in a that is well within of zero, but using both updated masses yields a positive proton separation energy consistent with proton-bound 150Tm ().
+The Er isotopic chain was measured alongside Tm during experiment 2 from 154Er to 149Er.
+During this second measurement, a discrepancy in the mass of 149gEr relative to the literature value was found that significantly affects the one-proton separation energy () of the 150Tm ground state.
+Together, the newly measured 150Tm mass and the AME2020 mass of 149gEr results in a that is well within of zero, but using both updated masses yields a positive proton separation energy consistent with proton-bound 150Tm ().
 As with 151Yb, for which the ISAC target was seen to preferentially deliver the spin 11/2 isomer over the ground state, 149Er, having two fewer protons, similarly appeared to favour the excited state.
@@
 The other method is to obtain the mass directly from the fit of the ground state within the tail of the isomer using the known peak shape.
-The first approach obtains the mass of 149gEr using the known isomeric excitation energy which was initially measured via a two -ray internal transition to the ground state proceeding through a 630.5 keV (M4) 111.0 keV (M1) cascade [58] This was later confirmed in subsequent experiments which yielded 630.3 keV 111.3 keV [59] and 630.52.6 keV 111.31.1 keV by [60].
+The first approach obtains the mass of 149gEr using the known isomeric excitation energy, which was initially measured via a two -ray internal transition to the ground state proceeding through a 630.5 keV (M4) 111.0 keV (M1) cascade [57] This was later confirmed in subsequent experiments which yielded 630.3 keV 111.3 keV [51] and 630.52.6 keV 111.31.1 keV by [58].
 The 149mEr isomer is established to lie 741.80.2 keV above the ground state according to the NUBASE 2020 evaluation [50].
@@
 However, the present dataset also enables a more direct measurement of the ground state mass by utilizing the lower abundance, overlapping TOF peak of the ground state of 149Er and employing a hyperEMG fitting procedure.
-Directly fitting the lower abundance 149gEr (as in Figure 1) with the hyperEMG function calibrated to the shape of the 149Dy peak yields a mass excess of -5358447 keV, an increase in mass of 15855 keV from the AME2020 value.
+Directly fitting the lower abundance 149gEr (as in Fig.
+1) with the hyperEMG function calibrated to the shape of the 149Dy peak yields a mass excess of -5358447 keV, an increase in mass of 15855 keV from the AME2020 value.
 The excitation energy of 149mEr can then be determined from the difference between our measurements of ground state and isomer to be 72653 keV.
@@
 The cause of our discrepancy in our measurement of this single mass is unknown, but may warrant a third measurement of the 149g,mEr isotope for confirmation.
-These new masses of the 149Er isomer and ground state lead us to adopt a revised one-proton separation energy of the ground state which directly impacts the location of the proton drip-line in Tm, since the proton separation energy of a nucleus depends on both the mass of the candidate proton emitter as well as on the mass of the daughter nucleus of the decay.
+These new masses of the 149Er isomer and ground state lead us to adopt a revised one-proton separation energy of the ground state, which directly impacts the location of the proton drip-line in Tm, since the proton separation energy of a nucleus depends on both the mass of the candidate proton emitter as well as on the mass of the daughter nucleus of the decay.
 Both the 149,150Tm masses as well as the revised 149Er value were thus needed for an unambiguous identification of the Tm drip-line, since the new Tm masses alone did not exclude the possibility that 150Tm could also be proton-unbound when using the 149Er value from the AME2020.
 These measurements confirm that 150Tm is indeed proton-bound, agreeing qualitatively with the Bayesian model averaging predictions of [30] that 149Tm is the first proton-unbound thulium isotope.
-Furthermore, the proton emission Q-value of 149Tm has been determined and can now be used as an input for decay rate calculations.
+Furthermore, the proton emission -value of 149Tm has been determined and can now be used as an input for decay rate calculations.
 149Tm: The First p-Unbound Isotope
 The proton emission energy can be determined from .
-Our direct measurements of 149gEr and 150gTm show 150gTm to be proton-bound, having a Q-value of -20299 keV.
+Our direct measurements of 149gEr and 150gTm show 150gTm to be proton-bound, having a -value of -20299 keV.
 Meanwhile, our measurement of 149gTm establishes it as the first proton-unbound isotope in the thulium chain.
-For it we obtain a proton emission Q-value of +37852 keV when calculated using the measured mass of 149gTm and the masses of 148Er and 1H from the AME2020.
+For it we obtain a proton emission -value of +37852 keV when calculated using the measured mass of 149gTm and the masses of 148Er and 1H from the AME2020.
 This decay energy is on the high side of the AME extrapolation of 250201 keV and can be used in estimations of the proton emission half-life for a given angular momentum transfer (e.g.
 [3, 4]).
-The small Q-value of the 149gTm decay allows for it to proceed to only one possible final state, the 0+ ground state of 148Er, so it is expected that J.
-Both of the above decay Q-values and their uncertainties take into account the expected unresolved isomers listed in NUBASE inferred from trends in neighbouring nuclei.
+The small -value of the 149gTm decay allows for it to proceed to only one possible final state, the 0+ ground state of 148Er, so it is expected that J.
+Both of the above decay -values and their uncertainties take into account the expected unresolved isomers listed in NUBASE inferred from trends in neighbouring nuclei.
 Evolution of the Neutron Shell Closure
-In addition to determining the location of the drip-line, the new masses of 149/150Tm allow for a comparison to predictions of the empirical shell gap of the shell closure defined by (see Figure 5).
+In addition to determining the location of the drip-line, the new masses of 149/150Tm allow for a comparison to predictions of the empirical shell gap of the shell closure defined by (see Fig.
+5).
 The models shown include UNEDF0 [60], UNEDF1 [26], HFB21 [61], Duflo-Zuker [62], FRDM [63], SkP [64], SV-min [27], and ETFSI-Q [65].
-In this region most models over-predict the shell gap of the N=82 isotone chain by 0.5-3 MeV with the exceptions of Duflo-Zuker under-predicting by nearly 1 MeV and HFB21 which lies within 500 keV of the isotones from to .
-The present work measures the of Tm at the N=82 shell closure and shows that it is, in fact, the first isotone for which depends on the binding energy of a proton-unbound nucleus (149Tm).
+In this region most models over predict the shell gap of the isotone chain by 0.5–3 MeV with the exceptions of Duflo-Zuker under predicting by nearly 1 MeV and HFB21 which lies within 500 keV of the isotones from to .
+The present work measures the of Tm at the shell closure and shows that it is, in fact, the first isotone for which depends on the binding energy of a proton-unbound nucleus (149Tm).
 These measurements indicate the persistence of the shell closure, which is consistent with the recent corresponding measurement of Yb [15].
-Furthermore, the empirical shell gap at Z=69 continues the trend seen in the more stable nuclei such as Z=63 and 67 of having a local maximum in at the odd-Z nuclei.
+Furthermore, the empirical shell gap at continues the trend seen in the more stable nuclei such as and 67 of having a local maximum in at the odd- nuclei.
 IV Conclusion and Outlook
@@
 Heavier exotic species pose challenges that arise from the abundance of isomers and isobaric molecules that must be identified and, therefore, must be treated with caution.
-The mass values presented here provide a new benchmark for the models used to predict the binding energies of unmeasured nuclei in the neutron-deficient region near N=82 and establish the Z=69 proton drip-line.
+The mass values presented here provide a new benchmark for the models used to predict the binding energies of unmeasured nuclei in the neutron-deficient region near and establish the proton drip-line.
 V Acknowledgements
@@
 05P16RGFN1 and No.
-05P19RGFN1; by the Hessian Ministry for Science and Art through the LOEWE Center HIC for FAIR; by the JLU and GSI Helmholtzzentrum für Schwerionenforschung under the JLU-GSI strategic Helmholtz partnership agreement.
-For the purpose of open access, the authors have applied a creative commons attribution (CC BY) licence to any author accepted manuscript version arising.
+05P19RGFN1; by the Hessian Ministry for Science and Art through the LOEWE Center HIC for FAIR; and by the JLU and GSI Helmholtzzentrum für Schwerionenforschung under the JLU-GSI strategic Helmholtz partnership agreement.
 | Species | Calibrant | Ion Mass Ratio | ME (keV) | ME (keV) |